# Swarming in the Wild: A Distributed Communication-less Lloyd-based Algorithm dealing with Uncertainties

### v1

@@ III-B The weighting function $\varphi_{i}(q)$
$\begin{split}\dot{\bar{p}}_{i}&=\begin{cases}-(\bar{p}_{i}-R(\pi/2-\varepsilon)e_{i})\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{3}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{4}\\ -(\bar{p}_{i}-e_{i})&\text{otherwise,}\end{cases}\\ \bar{p}_{i}&=e_{i}\,\,\,\text{if}\,\,\,\|p_{i}-\bar{c}_{\mathcal{A}_{i}}\|>\|p_{i}-c_{\mathcal{A}_{i}}\|\,\land\,\bar{p}_{i}=R(\pi/2-\varepsilon)e_{i}.\\ \end{split}$ (12)
The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the $i$–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
Corollary 1 (Convergence in cluttered environment and with flocking constraints).
Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$ and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other. The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. By imposing the control law $\dot{p}_{i}(\mathcal{F}_{i})$ in (4),  (10), (11), and (12) to compute the centroid position in (5), by properly selecting $d_{1},d_{2},d_{3},d_{4}$, and $\beta^{D}_{i}$, we have that $\lim_{t\rightarrow\infty}\|p_{i}(t)-e_{i}\|<r_{s,i},\forall i=1,\dots,N.$
Figure 4: We depict in cyan the robots and in grey the static obstacles. The crosses represent the final goal position while the dashed lines the proximity constraint topology. In the figure we show two possible deadlock conditions: case 1. is due to the small inter-space between obstacles. case 2. is due to the too short $\Gamma_{ij}$ value (or too big $\delta^{o}$ value).
@@
With respect to the proof of convergence with no obstacles and no proximity maintenance constraints, which are provided in [2], we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
Notice that the addition of static obstacles is equivalent to non-cooperative static robots. In order to not affect the proof of convergence obtained in [2], two requirements need to be considered. 1. We require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4). This is a consequence of the results in [2]. In fact, static obstacles can be thought as robots that reached their goal and have a value of $\beta_{i}\rightarrow 0$. However, from [2], by design this is not possible if the robots are at a distance from their goal smaller than $d_{2}$. Since $d_{2}>2\delta$ and $\|e_{i}-e_{j}\|>\Delta_{ij}$, then the condition $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$ is equivalent. 2. We require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4). By imposing these additional conditions, the guarantees of convergence in [2] holds. ∎
@@ III-C Dealing with dynamics and uncertainties
The real robot deployment requires to consider a set of additional issues, such as 1. More complex dynamics of the robots must be applied, i.e., $\dot{p}_{i}=u_{i}$ (Assumption 4) cannot be directly applied. 2. Tracking error, which can be due to external disturbances or model mismatches, generating a difference between the desired and the actual velocity, 3. Measurement uncertainties, due to the limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each $i$–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell. By doing that, we can account for the robot dynamics without violating the safety and proximity maintenance constraints. Hence, each robot $i$ solves the following optimization problem (for the sake of clarity we discard the $i$ indexing),
@@
$\,{x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U},\,\,\,\forall k\in(0,N_{t})$ (13c)
$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t})$ (13d)
@@
Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy. In fact, it can be achieved by adapting the value of $\beta_{i}$ in Equation (11) in an online fashion. An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [10]. However, this method cannot be applied in our case, since, by following this approach, the robot may fall out of its cell due to tracking error, and once it happens, both the cell and the centroid position are no longer defined. Conversely, the value of $\beta_{i}$ defines the aggressiveness of the $i$–th robot. A smaller value corresponds to a centroid position closer to the goal (and hence closer to the boundaries of the $\mathcal{F}_{i}$ cell, namely $\partial\mathcal{F}_{i}$), while larger values of $\beta_{i}$ push the centroid position as far as possible from the boundaries of the cell $\partial\mathcal{F}_{i}$. Thus, managing errors and uncertainties can be achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell. This distance must be chosen in accordance with the maximum positional tracking error and the maximum error in the neighbors’ position estimates. More formally, it can be done by imposing a lower bound to $\beta_{i}$ in (11) equal to
@@
where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point on the contour of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties. For the sake of clarity, we depicted three scenarios in Figure 5. Figure 5a considers the tracking error; in this case the value of $c_{\mathcal{F}_{i}}$ is controlled to stay at least at a distance $d^{t}_{u}$ from the boundary $\partial\mathcal{F}_{i}$. We indicate with $\mathcal{Z}_{i}$ the region set where the $i$–th robot is allowed to go in the next time step. Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$. To be conservative, the value of $d_{u}^{t}$ should be determined experimentally as the maximum tracking error, and it can be set as a function of the velocity. In Figure 5b, we show the case where we have measurement uncertainties. We illustrate both the $i$–th and the $j$–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$. In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety. By relying on geometric reasoning, depending on the measurement uncertainty, when $\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|<k\sqrt{\lambda^{\max}}+\delta$, we can tune the parameter $d_{u}^{m}$ as
$d_{u}^{m}\geq\left|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right|+\delta,$ (14)
@@
From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (14) in the worst-case expected scenario.
In real-world equipments the uncertainties associated with the measurements are far from being constant values. Hence, the user may instead set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means. One method to do this, is to use the restraining method from [31]. This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
On the other hand, to preserve proximity maintenance, if $\Gamma_{ij}-\|p_{i}-p_{j}\|\approx k\sqrt{\lambda^{\max}}$, we need to set $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$. Since this value may be quite high in practice, it is sufficient to impose that if the mean value of the neighbor position is out of the proximity area, then we can re-project its position to the closest point inside the proximity area. In this way, we can effectively maintain the proximity with the neighbors.
@@
By relying on the adaptive solution, while safety and proximity maintenance have been taken into account, we lost the convergence guarantees. In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$ and $\epsilon_{o}$. In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $d_{u}<d^{\min}_{o}-\sqrt{2}(\delta_{i}+\delta^{o})$. Notice that, if $d^{\min}_{o}<2\sqrt{2}(\delta_{i}+\delta^{o})$, the value of $\epsilon_{o}$ also plays a role. Hence, in this case, either the value of $d_{u}$ should be selected in a less conservative way, or we need to select $\epsilon_{o}=1$. A situation where the cell is too small to allow the centroid to be at the required distance from the boundary of the cell can also occur. However, we claim that this is not an issue for safety nor for proximity maintenance, since in these cases the robot velocity approaches zero. Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots which are close to the maximum allowed distance $\Gamma_{ij}$.
@@ IV Simulation results
In this section we show through simulations, the scalability, repeatability and robustness of our approach, and we provide a comparison with a state-of-the-art algorithm. In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with our proposed algorithm (refer to Table I). Despite [7] performing well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios. In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning. This is a common issue in the state-of-the-art algorithms, which is overcome by our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings [2]. We perform the comparison in the MRS simulator [32] in the same environment and same model dynamics for both algorithms. We simulate three scenarios with different number of robots and with different increasingly complex obstacle configurations. For each scenario, we perform $10$ runs for statistical evidence. The metrics that we took into account are:
  • The time required for all the robots to travel $42$ meters in the goal direction;
  • The success rate in terms of mutual collision avoidance and obstacle avoidance.
In Table II, we report the results obtained from PACNav [7] and from our approach. We considered the following scenario: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we selected the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m). The adjacency matrix which defines the proximity maintenance constraints were selected as,
$A_{d}=\begin{bmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{bmatrix},$
and
$A_{d}=\begin{bmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{bmatrix}.$
For our approach we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$. The parameter selection for PACNav [7] has been done by trial and error, on the basis of the environment and the number of robots. Besides that, as seen in from Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
@@
Scenario avg. time (s) SR avg. time (s) SR avg. time (s) SR
1 55.01 1.00 39.96 1.00 24.23 1.00
2 - 0.00 45.71 1.00 29.53 1.00
3 - 0.00 120.33 1.00 55.87 1.00
Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in $3$ different scenarios, comparing PACNav [7] with our method in two different parameter configurations.

### v2

@@ III-B The weighting function $\varphi_{i}(q)$
[…]
The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the ii–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
Proposition 1 (Convergence in cluttered environment and with flocking constraints).
Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$, where $\delta=\delta_{i},\forall i=1,\dots,N$, and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other. The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}\geq\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. By imposing the control law $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), (10), (11), and (12) to compute the centroid position in (5), by properly selecting $d_{1},d_{2},d_{3},d_{4}$, and $\beta^{D}_{i}$, we have that $\lim_{t\rightarrow\infty}\|p_{i}(t)-e_{i}\|<r_{s,i},\forall i=1,\dots,N.$
Figure 4: The figure shows two possible deadlock conditions: case 1. is due to the small inter-space between obstacles. case 2. is due to the too short $\Gamma_{ij}$ value (or too big $\delta^{o}$ value). We depict in cyan the robots and in grey the static obstacles. The crosses represent the final goal positions while the dashed lines represent the proximity constraint topology.
@@
The proof provided in [2] does not consider presence of static obstacles nor proximity constraints. By including both of them, we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
Notice that the addition of static obstacles is equivalent to non-cooperative static robots. In order to not affect the proof of convergence obtained in [2], two requirements need to be considered. First, we require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4). This is a consequence of the results in [2]. In fact, static obstacles can be thought as robots that reached their goal and have a value of $\beta_{i}\rightarrow 0$. However, from [2], by design this is not possible if the robots are at a distance from their goal smaller than $d_{2}$. Since $d_{2}>2\delta$ and $\|e_{i}-e_{j}\|>\Delta_{ij}$, then the condition $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$ is equivalent. Second, we require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4). By imposing these additional conditions, the guarantees of convergence in [2] holds. ∎
@@ III-C Dealing with dynamics and uncertainties
The real robot deployment requires to consider a set of additional issues. First, more complex dynamics of the robots must be applied as $\dot{p}_{i}=u_{i}$ is not directly applicable. Second, tracking error, which can be caused by external disturbances or model mismatches, generating a difference between the desired and the actual velocity, has to be taken in to account. Finally, we have to consider measurement uncertainties arising from limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each ii–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell. By doing that, we can account for the robot dynamics without violating the safety and proximity maintenance constraints. Hence, each robot $i$ solves the following optimization problem (for the sake of clarity we discard the $i$ indexing),
@@
$\,{x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U},\,\,\,\forall k\in(0,N_{t}),$ (13c)
$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t}),$ (13d)
@@
Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy. In fact, it can be achieved by adapting the value of $\beta_{i}$ in (11) in an online fashion. An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [10]. However, this method cannot be applied in our case, since, by following this approach, the robot may fall out of its cell due to tracking error, and once it happens, both the cell and the centroid position are no longer defined. Conversely, the value of $\beta_{i}$ defines the aggressiveness of the ii–th robot. A smaller value corresponds to a centroid position closer to the goal (and hence closer to the boundaries of the $\mathcal{F}_{i}$ cell, namely $\partial\mathcal{F}_{i}$), while larger values of $\beta_{i}$ push the centroid position as far as possible from the boundaries of the cell $\partial\mathcal{F}_{i}$. Thus, managing errors and uncertainties is achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell. This distance must be chosen in accordance with the maximum positional tracking error and the maximum error in the neighbors’ position estimates. More formally, it can be done by imposing a lower bound to $\beta_{i}$ in (11) equal to
@@
where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point on the contour of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties. For the sake of clarity, we depicted three scenarios in Figure 5. Figure 5a considers the tracking error; in this case the value of $c_{\mathcal{F}_{i}}$ is controlled to stay at least at a distance $d^{t}_{u}$ from the boundary $\partial\mathcal{F}_{i}$. We indicate with $\mathcal{Z}_{i}$ the region set where the ii–th robot is allowed to go in the next time step. Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$. To be conservative, the value of $d_{u}^{t}$ should be determined experimentally as the maximum tracking error, and it can be set as a function of the velocity. In Figure 5b, we show the case where we have measurement uncertainties. We illustrate both the ii–th and the jj–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$. In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety. By relying on geometric reasoning, depending on the measurement uncertainty, when $\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|<k\sqrt{\lambda^{\max}}+\delta$, we can tune the parameter $d_{u}^{m}$ as
$d_{u}^{m}\geq\left\|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right\|+\delta,$ (14)
@@
From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (14) in the worst-case expected scenario. In real-world equipments the uncertainties associated with the measurements are far from being constant values. Hence, the user can set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means. One method to implement this, is to use the restraining method from [31]. This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
On the other hand, to preserve proximity maintenance, if $\Gamma_{ij}-\|p_{i}-p_{j}\|\approx k\sqrt{\lambda^{\max}}$, we need to set $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$. Since this value may be quite high in practice, it is sufficient to impose that if the mean value of the neighbor position is out of the proximity area, then we can re-project its position to the closest point inside the proximity area. In this way, the proximity to the neighbors is effectively maintained.
@@
By relying on the adaptive solution, while safety and proximity maintenance have been taken into account, we lost the convergence guarantees. In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$. In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $0\leq d_{u}<\frac{d^{\min}_{o}}{2}-(\delta_{i}+\delta^{o})$. A situation where the cell is too small to allow the centroid to be at the required distance from the boundary of the cell can also occur. However, we claim that this is not an issue for safety nor for proximity maintenance, since in these cases the robot velocity approaches zero. Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots that are close to the maximum allowed distance $\Gamma_{ij}$.
@@ IV Simulation results
In this section we show through simulations, the scalability, repeatability and robustness of our approach, and we provide a comparison with a state-of-the-art algorithm. In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with the proposed algorithm (refer to Table I). Notice that comparing with algorithm such as [5, 4, 6] is out of scope of this paper. Our method does not perform better than algorithms that can share information, such as future trajectories, among the robots. Despite [7] performs well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios. In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning. This is a common issue in the state-of-the-art algorithms, which is overcome in our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings. We perform the comparison in the MRS simulator [32] in the same environment and same model dynamics for both algorithms. We simulate three scenarios with different number of robots and with an increasingly complex obstacle configurations. For each scenario, we perform $10$ runs for statistical evidence. The metrics that we took into account are the time required for all the robots to travel $42$ meters in the goal direction and the success rate in terms of mutual collision avoidance and obstacle avoidance. In Table II, we report the results obtained from PACNav [7] and from our approach. We considered the following scenario: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we selected the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m). The adjacency matrix which defines the proximity maintenance constraints were selected as,
$A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right],\text{ and}\quad A_{d}=\left[\begin{smallmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{smallmatrix}\right].$
For our approach we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$. The parameter selection for PACNav [7] has been done by trial and error, on the basis of the environment and the number of robots. Besides that, as seen in Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
@@
Scenario N avg. time (s) SR avg. time (s) SR avg. time (s) SR
1 3 55.01 1.00 39.96 1.00 24.23 1.00
2 9 - 0.00 45.71 1.00 29.53 1.00
3 9 - 0.00 120.33 1.00 55.87 1.00
Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in 33 different scenarios, comparing PACNav [7] with our method in two different parameter configurations.
For further simulation results, please refer to the attached multimedia material, where we include simulations with up to $16$ UAVs in crowded conditions. Notice that even if the sufficient condition in Proposition 1 is not respected, the robots can still safely reach the goal in environments with $56\%$ of free space, as shown in multimedia material.